# Fast Radio Bursts and Electromagnetic Transition Radiation on Gravitational Shockwaves

### v1

@@ Abstract
Introduction. Fast radio bursts are bright pulses of radio frequencies between 400 MHz and 8 GHz. Although the origins of FRBs are still unknown, there has been significant progress in understanding their possible physics [1] in recent years. Some, and probably all, of these bursts are generated by the release of energy from a magnetar [2]. This process involves two mechanisms: the primary perturbation of the magnetar’s magnetosphere and the subsequent conversion of the perturbation energy into a radio burst. Observational data make it possible to significantly restrict the class of models describing the second mechanism — the emission of radio waves. One of the most successful models is the synchrotron maser emission [3, 4]. This process occurs at a noticeable distance from the magnetar and is initiated by the primary short intense outburst of energy from the magnetosphere [5, 6]. However, the primary perturbation of the magnetosphere can be caused by various reasons. If a blast satisfies some basic criteria [7], it, regardless of its initial nature, can potentially generate an observed FRB. Therefore, it is important to study all the fundamental mechanisms, the FRB ‘engines’, that lead to the high-energy outburst from magnetar.
@@
The engines can be divided into two main classes: those caused by internal processes of the magnetar, and those caused by external impact. Among external impacts, mainly collisions are considered: with asteroids [8], another neutron star [9], black hole [10] or other massive compact objects [11]. Interactions without collision are also considered [12]. The aim of the present work is to demonstrate a novel ‘collisionless’ mechanism of external impact leading to a short powerful outburst of radiation from the magnetar. This is a transition radiation of the magnetar dipole moment under the action of a gravitational shockwave (GSW).
@@
The transition radiation has been described in 1945 by Ginzburg and Frank [13] as a universal phenomenon which occurs when a compact charged source moves from one region of space to another with different properties. In the original version, this was the radiation of a uniformly moving point charge crossing the boundary of two media with different permittivities. An alternative scenario is the incidence of a permittivity wave (associated, for example, with a density wave) on a charge at rest [14]. Obviously, the solution of Maxwell equations is affected not only by a change in permittivity, but also by a change in the spacetime geometry. Therefore, Ginzburg and Tsytovich subsequently considered the transition radiation that occurs when a weak gravitational wave falls on a compact source of an electromagnetic field [15]. In particular, they studied the scattering of a gravitational wave by a magnetar and the formation of electromagnetic transition radiation as well as Alfven waves in this process. However, for the weak gravitational wave, the resulting radiation also turned out to be too weak to lead to noticeable observable effects. Alternative considerations of energy transfer between gravitational and electromagnetic waves in an external stationary electromagnetic field have repeatedly been encountered in the literature, e.g., the Gertsenshtein effect [16], see also [17].
@@
A natural way to enhance the power of the transition radiation on a gravitational wave is to consider strong gravitational waves, and gravitational shockwaves, in particular. GSW are solutions to the Einstein equations which describe short gravitational pulses with rapid oscillations of the Riemann curvature on the shockwave fronts. In general, the Einstein and Maxwell equations then become essentially nonlinear in the metric, which leads to significant computational difficulties. In this work we study plane-fronted shockwaves which propagate in flat space-time or in a space-times where gravitational effects are weak with respect to the gravitational shocks. Such gravitational shockwaves are known to be created by the motion of ultrarelativistic sources and have a unique feature: their Einstein tensor is linear in the metric. These GSW include the subclass of Kundt spacetimes [18].
@@
In our approximation, test particles before and after the shockwave pulse move along geodesics in the Minkowsky space-time. After the shockwave, mutual orientations of 4-velocities of the particles change, as a manifestation of the gravitational memory effect. Transformations of the 4-velocities on the shockwave front look as coordinate transformations determined by the profile of the GSW. These transformations are also known as the Penrose supertranslations [19]. Analogous effects hold for the transition radiation and play the crucial role in the subsequent analysis.
The main results of the present work are two-fold. First, we demonstrate that the gravitational memory effects on electromagnetic perturbations caused by the considered GSW have a very simple form. Similar to trajectories of particles the transition radiation ‘remembers’ the profile of a GSW but does not depend on the complexity of the gravitational ‘signal’. This fact allows one to overcome computational difficulties relatively easily. Second, on this base, we study perturbations of the magnetar’s magnetic field caused by GSW created by different ultrarelativistic objects which move near a magnetar. We analyze the angular distribution of the transition radiation flux, the duration of the pulses and show that they are consistent with properties of the FRB engines. Our suggestion for the novel FRB engine mechanism is illustrated on Fig. 1.
Gravitational memory of trajectories after shockwaves. We start with necessary definitions and illustration of the gravitational memory effect for geodesics on the considered class of plane-fronted shockwave geometries. For a shock moving along the $x$ direction, from the left to the right, the space-time metric is chosen to be
$ds^{2}=-dvdu-H(v,u,y)du^{2}+dy_{i}^{2}~{}.$ (1)
@@
$G_{uu}=\frac{1}{2}\partial^{2}_{i}H~{},~{}G_{uv}=\partial_{v}\partial_{i}H~{},~{}G_{ij}=2\delta_{ij}\partial_{v}^{2}H~{}.$ (3)
@@
$T_{\mu\nu}=\chi(u)\left(\sigma l_{\mu}l_{\nu}+j_{i}(l_{\mu}e_{\nu}^{i}+l_{\nu}e_{\mu}^{i})+p\delta_{ij}e_{\mu}^{i}e_{\nu}^{j}\right)~{},$ (4)
where $\sigma=f_{,ii}/\kappa$, $j_{i}=-2f_{,vi}/\kappa$, $p=4f_{,vv}/\kappa$, $\kappa=16\pi G$. Vectors $e^{i}_{\mu}=\delta^{i}_{\mu}$ are tangent to null hypersurfaces of constant $u$. Concrete examples of the profile functions are given below for ultrarelativistic particles and cosmic strings.
@@
$u^{\mu}_{1}=-\frac{1}{2}l^{\mu}f~{},~{}u^{\mu}_{2}=-\frac{1}{2}(f_{,\nu}\bar{u}^{\nu})-\frac{1}{2}l^{\mu}f~{}.$ (6)
@@
$\delta x^{\mu}=-\zeta^{\mu}~{},~{}u^{\mu}\simeq\bar{u}^{\mu}+{\cal L}_{\zeta}\bar{u}^{\mu}~{},~{}u\simeq\delta~{},$ (7)
@@
$\zeta^{\mu}(x)\simeq f\delta^{\mu}_{v}+\frac{u}{2}\eta^{\mu a}f_{,a}~{}.$ (8)
Here ${\cal L}_{\zeta}$ is the Lie derivative, vector $\zeta^{\mu}$ has been introduced in [20]. To derive (8) we took into account that, in accord with the chosen normalization, $\bar{\chi}(u)=1$ at $u>\delta$. Coordinate transformations at $u=0$ are well-known Penrose supertranslations [19]. Also they make an infinite dimensional group of isometries of the surface $u=0$, known as Carroll transformations [21, 22, 23, 24].
@@
Gravitational memory of EM fields. Our primary interest is to determine perturbations of classical EM fields caused by shockwaves by using the gravitational memory effect. Consider Maxwell equations on shockwave spacetime (1) and in Minkowsky spacetime
$\partial_{\mu}F^{\mu\nu}=j^{\nu}~{},~{}\partial_{\mu}\bar{F}^{\mu\nu}=\bar{j}^{\nu}~{},$ (9)
@@
$J^{\nu}_{g}(F,f)=\partial_{\mu}\left[fl^{\alpha}(l^{\nu}\eta^{\mu\beta}-l^{\mu}\eta^{\nu\beta})F_{\alpha\beta})\right]~{}.$ (11)
It is clear that $\partial_{\mu}J^{\mu}=\partial_{\mu}(\chi(u)J^{\mu}_{g})=\partial_{\mu}J^{\mu}_{g}=0$ and $l_{\mu}J^{\mu}_{g}=0$. In the first approximation one can put $J^{\nu}_{g}(F,f)\simeq J^{\nu}_{g}(\bar{F},f)$. After some algebra one can write
$J^{\nu}_{g}(\bar{F},f)=-2fl_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-fl^{\nu}l_{\alpha}\partial_{\lambda}\bar{F}^{\alpha\lambda}~{}.$ (12)
@@
$p_{(\mu\nu}l_{\rho)}=0~{},~{}p_{(\mu\nu,\rho)}=q_{(\mu\nu}l_{\rho)}~{},$ (15)
$l_{\mu}i_{1}^{\mu}=0~{},~{}\partial_{\mu}i_{1}^{\mu}=l_{\nu}i_{2}^{\nu}~{},~{}\partial_{\mu}i_{2}^{\mu}=0~{},$ (16)
where symbol $(\mu\nu...\rho)$ denotes complete symmetrization of the indices. Indices are risen and lowered with the flat metric. Substitution of (13), (14) to Maxwell equations (10) yields:
$l_{\mu}p^{\mu\nu}$ $=0~{},$ (17)
$\partial_{\mu}p^{\mu\nu}-l_{\mu}q^{\mu\nu}$ $=i_{1}^{\nu}+J_{g}^{\nu}~{},$ (18)
$\partial_{\mu}q^{\mu\nu}$ $=i_{2}^{\nu}~{}.$ (19)
@@
$p_{\mu\nu}=\sigma_{i}(l_{\mu}e^{i}_{\nu}-l_{\nu}e^{i}_{\mu})~{},$ (20)
$q_{\mu\nu}=e^{i}_{\mu}\sigma_{i,\nu}-e^{i}_{\nu}\sigma_{i,\mu}+\rho_{i}(l_{\mu}e^{i}_{\nu}-l_{\nu}e^{i}_{\mu})~{},$ (21)
@@
$l_{\mu}q^{\mu\nu}=l_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-{\cal J}^{\nu}~{},$ (23)
${\cal J}^{\nu}=-\frac{1}{2}(i_{1}^{\nu}+fl^{\nu}l_{\alpha}{\bar{j}}^{\alpha}+\sigma_{i,i}l^{\nu})~{}.$ (24)
@@
$l_{\mu}\hat{F}^{\mu\nu}\simeq\bar{\chi}(u)\left(l_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-{\cal J}^{\nu}\right)~{}.$ (25)
@@
Formulation of the problem. We are interested in perturbation $\hat{F}_{\mu\nu}\equiv F_{\mu\nu}-\bar{F}_{\mu\nu}$ in the region $u>\delta$, where both $F_{\mu\nu}$ and $\bar{F}_{\mu\nu}$ are solutions to Maxwell equations in Minkowsky space time. We require that in this region $j^{\mu}={\bar{j}}^{\mu}$. Therefore, $\hat{F}_{\mu\nu}$ obeys homogeneous Maxwell equations. Our results allow to formulate the following characteristic Cauchy problem:
$\partial^{\mu}\hat{F}_{\mu\nu}=0~{},~{}u>\delta~{},$ (26)
$l^{\mu}\hat{F}_{\mu\nu}=l^{\mu}{\cal L}_{\zeta}\bar{F}_{\mu\nu}~{},~{}u=\delta~{}~{},$ (27)
@@
$\hat{F}_{uv}=-\frac{1}{2}\int_{-\infty}^{v}\partial_{i}\hat{F}_{vi}~{}dv^{\prime}~{}.$ (28)
@@
$\Box\hat{F}_{vi}=0~{},~{}u>0~{},$ (30)
$\hat{F}_{vi}={\cal L}_{\zeta}\bar{F}_{vi}=\partial_{v}(f\bar{F}_{vi})~{},~{}u=0~{}.$ (31)

### v2

# 141980, Dubna, Russia

@@ 1 Introduction
Fast radio bursts are bright pulses of radio frequencies between 400 MHz and 8 GHz. Although the origins of FRBs are still unknown, there has been significant progress in understanding their possible physics Petroff:2021wug in recent years. Some, and probably all, of these bursts are generated by the release of energy from a magnetar Zhang:2022uzl . This process involves two mechanisms: the primary perturbation of the magnetar’s magnetosphere and the subsequent conversion of the perturbation energy into a radio burst. Observational data make it possible to significantly restrict the class of models describing the second mechanism — the emission of radio waves. One of the most successful models is the synchrotron maser emission Lyubarsky:2014jta ; Beloborodov:2017juh . This process occurs at a noticeable distance from the magnetar and is initiated by the primary short intense outburst of energy from the magnetosphere Metzger:2019una ; Beloborodov:2019wex . However, the primary perturbation of the magnetosphere can be caused by various reasons. If a blast satisfies some basic criteria Margalit:2019tph , it, regardless of its initial nature, can potentially generate an observed FRB. Therefore, it is important to study all the fundamental mechanisms, the FRB ‘engines’, that lead to the high-energy outburst from magnetar.
@@
The engines can be divided into two main classes: those caused by internal processes of the magnetar, and those caused by external impact. Among external impacts, mainly collisions are considered: with asteroids Dai:2016qem , another neutron star Moroianu:2022ccs , black hole Kim:2024fuy or other massive compact objects Liu:2020zyj . Interactions without collision are also considered Zhang:2020eou . The aim of the present work is to demonstrate a novel ‘collisionless’ mechanism of external impact leading to a short powerful outburst of radiation from the magnetar. This is a transition radiation of the magnetar dipole moment under the action of a gravitational shockwave (GSW).
@@
The transition radiation has been described in 1945 by Ginzburg and Frank Ginzburg:1945zz as a universal phenomenon which occurs when a compact charged source moves from one region of space to another with different properties. In the original version, this was the radiation of a uniformly moving point charge crossing the boundary of two media with different permittivities. An alternative scenario is the incidence of a permittivity wave (associated, for example, with a density wave) on a charge at rest Ginzburg:1979wi . Obviously, the solution of Maxwell equations is affected not only by a change in permittivity, but also by a change in the spacetime geometry. Therefore, Ginzburg and Tsytovich subsequently considered the transition radiation that occurs when a weak gravitational wave falls on a compact source of an electromagnetic field Ginzburg:1975 . In particular, they studied the scattering of a gravitational wave by a magnetar and the formation of electromagnetic transition radiation as well as Alfven waves in this process. However, for the weak gravitational wave, the resulting radiation also turned out to be too weak to lead to noticeable observable effects. Alternative considerations of energy transfer between gravitational and electromagnetic waves in an external stationary electromagnetic field have repeatedly been encountered in the literature, e.g., the Gertsenshtein effect Gertsenshtein:1962 , see also Wen:2014wxa .
@@
A natural way to enhance the power of the transition radiation on a gravitational wave is to consider strong gravitational waves, and gravitational shockwaves, in particular. GSW are solutions to the Einstein equations which describe short gravitational pulses with rapid oscillations of the Riemann curvature on the shockwave fronts. In general, the Einstein and Maxwell equations then become essentially nonlinear in the metric, which leads to significant computational difficulties. In this work we study plane-fronted shockwaves which propagate in flat space-time or in a space-times where gravitational effects are weak with respect to the gravitational shocks. Such gravitational shockwaves are known to be created by the motion of ultrarelativistic sources and have a unique feature: their Einstein tensor is linear in the metric. These GSW include the subclass of Kundt spacetimes Kundt:1961 .
In our approximation, test particles before and after the shockwave pulse move along geodesics in the Minkowsky space-time. After the shockwave, mutual orientations of 4-velocities of the particles change, as a manifestation of the gravitational memory effect. Transformations of the 4-velocities on the shockwave front look as coordinate transformations determined by the profile of the GSW. These transformations are also known as the Penrose supertranslations Penrose:1972xrn . Analogous effects hold for the transition radiation and play the crucial role in the subsequent analysis.
The main results of the present work are two-fold. First, we demonstrate that the gravitational memory effects on electromagnetic perturbations caused by the considered GSW have a very simple form. Similar to trajectories of particles the transition radiation ‘remembers’ the profile of a GSW but does not depend on the complexity of the gravitational ‘signal’. This fact allows one to overcome computational difficulties relatively easily. Second, on this base, we study perturbations of the magnetar’s magnetic field caused by GSW created by different ultrarelativistic objects which move near a magnetar. We analyze the angular distribution of the transition radiation flux, the duration of the pulses and show that they are consistent with properties of the FRB engines. Our suggestion for the novel FRB engine mechanism is illustrated on figure 1.
@@
2 Gravitational memory of trajectories after shockwaves
We start with necessary definitions and illustration of the gravitational memory effect for geodesics on the considered class of plane-fronted shockwave geometries. For a shock moving along the $x$ direction, from the left to the right, the space-time metric is chosen to be
$ds^{2}=-dvdu-H(v,u,y)du^{2}+dy_{i}^{2}\,.$ (1)
@@ 2 Gravitational memory of trajectories after shockwaves
$G_{uu}=\frac{1}{2}\partial^{2}_{i}H\,,\quad G_{ui}=\partial_{v}\partial_{i}H\,,\quad G_{ij}=2\delta_{ij}\partial_{v}^{2}H\,.$ (3)
@@
$T_{\mu\nu}=\chi(u)\left(\sigma l_{\mu}l_{\nu}+j_{i}(l_{\mu}e_{\nu}^{i}+l_{\nu}e_{\mu}^{i})+p\delta_{ij}e_{\mu}^{i}e_{\nu}^{j}\right)\,,$ (4)
where
$\sigma=f_{,ii}/\kappa\,,\quad j_{i}=-2f_{,vi}/\kappa\,,\quad p=4f_{,vv}/\kappa\,,\quad\kappa=16\pi G\,.$ (5)
Vectors $e^{i}_{\mu}=\delta^{i}_{\mu}$ are tangent to null hypersurfaces of constant $u$. Concrete examples of the profile functions are given below for ultrarelativistic particles and cosmic strings.
@@
$u^{\mu}_{1}=-\frac{1}{2}l^{\mu}f\,,\quad u^{\mu}_{2}=-\frac{1}{2}l^{\mu}(f_{,\nu}\bar{u}^{\nu})-\frac{1}{2}f^{,\mu}\,.$ (7)
@@
$\delta x^{\mu}=-\zeta^{\mu}\,,\quad u^{\mu}\simeq\bar{u}^{\mu}+{\cal L}_{\zeta}\bar{u}^{\mu}\,,\quad u\simeq\delta\,,$ (8)
@@
$\zeta^{\mu}(x)\simeq f\delta^{\mu}_{v}+\frac{u}{2}\eta^{\mu a}f_{,a}\,.$ (9)
Here ${\cal L}_{\zeta}$ is the Lie derivative, vector $\zeta^{\mu}$ has been introduced in Blau:2015nee . To derive (9) we took into account that, in accord with the chosen normalization, $\bar{\chi}(u)=1$ at $u>\delta$. Coordinate transformations at $u=0$ are well-known Penrose supertranslations Penrose:1972xrn . Also they make an infinite dimensional group of isometries of the surface $u=0$, known as Carroll transformations Duval:2014lpa ; Duval:2014uoa ; Ciambelli:2023tzb ; Ciambelli:2023xqk .
@@
3 Gravitational memory of EM fields
Our primary interest is to determine perturbations of classical EM fields caused by shockwaves by using the gravitational memory effect. Consider Maxwell equations on shockwave spacetime (1) and in Minkowsky spacetime
$\partial_{\mu}F^{\mu\nu}=j^{\nu}\,,\quad\partial_{\mu}\bar{F}^{\mu\nu}=\bar{j}^{\nu}\,,$ (10)
@@ 3 Gravitational memory of EM fields
$J^{\nu}_{g}(F,f)=\partial_{\mu}\left[fl^{\alpha}(l^{\nu}\eta^{\mu\beta}-l^{\mu}\eta^{\nu\beta})F_{\alpha\beta})\right]\,.$ (12)
It is clear that
$\partial_{\mu}J^{\mu}=\partial_{\mu}(\chi(u)J^{\mu}_{g})=\partial_{\mu}J^{\mu}_{g}=0\,,\quad l_{\mu}J^{\mu}_{g}=0\,.$ (13)
In the first approximation one can put $J^{\nu}_{g}(F,f)\simeq J^{\nu}_{g}(\bar{F},f)$. After some algebra one can write
$J^{\nu}_{g}(\bar{F},f)=-2fl_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-fl^{\nu}l_{\alpha}\partial_{\lambda}\bar{F}^{\alpha\lambda}\,.$ (14)
@@
$p_{(\mu\nu}l_{\rho)}=0\,,\quad p_{(\mu\nu,\rho)}=q_{(\mu\nu}l_{\rho)}\,,$ (17)
$l_{\mu}i_{1}^{\mu}=0\,,\quad\partial_{\mu}i_{1}^{\mu}=l_{\nu}i_{2}^{\nu}\,,\quad\partial_{\mu}i_{2}^{\mu}=0\,,$ (18)
where symbol $(\mu\nu\ldots\rho)$ denotes complete symmetrization of the indices. Indices are risen and lowered with the flat metric. Substitution of (15), (16) to Maxwell equations (11) yields:
$l_{\mu}p^{\mu\nu}$ =0=0\,,= 0 , (19)
$\partial_{\mu}p^{\mu\nu}-l_{\mu}q^{\mu\nu}$ $=i_{1}^{\nu}+J_{g}^{\nu}\,,$ (20)
$\partial_{\mu}q^{\mu\nu}$ $=i_{2}^{\nu}\,.$ (21)
@@
$p_{\mu\nu}=\sigma_{i}(l_{\mu}e^{i}_{\nu}-l_{\nu}e^{i}_{\mu})\,,$ (22)
$q_{\mu\nu}=e^{i}_{\mu}\sigma_{i,\nu}-e^{i}_{\nu}\sigma_{i,\mu}+\rho_{i}(l_{\mu}e^{i}_{\nu}-l_{\nu}e^{i}_{\mu})\,,$ (23)
@@
$l_{\mu}q^{\mu\nu}=l_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-{\cal J}^{\nu}\,,$ (25)
${\cal J}^{\nu}=-\frac{1}{2}(i_{1}^{\nu}+fl^{\nu}l_{\alpha}{\bar{j}}^{\alpha}+\sigma_{i,i}l^{\nu})\,.$ (26)
@@
$l_{\mu}\hat{F}^{\mu\nu}\simeq\bar{\chi}(u)\left(l_{\mu}{\cal L}_{\zeta}\bar{F}^{\mu\nu}-{\cal J}^{\nu}\right)\,.$ (27)
@@
4 Formulation of the problem
We are interested in perturbation $\hat{F}_{\mu\nu}\equiv F_{\mu\nu}-\bar{F}_{\mu\nu}$ in the region $u>\delta$, where both $F_{\mu\nu}$ and $\bar{F}_{\mu\nu}$ are solutions to Maxwell equations in Minkowsky space time. We require that in this region $j^{\mu}={\bar{j}}^{\mu}$. Therefore, $\hat{F}_{\mu\nu}$ obeys homogeneous Maxwell equations. Our results allow to formulate the following characteristic Cauchy problem:
$\partial^{\mu}\hat{F}_{\mu\nu}=0\,,\quad u>\delta\,,$ (28)
$l^{\mu}\hat{F}_{\mu\nu}=l^{\mu}{\cal L}_{\zeta}\bar{F}_{\mu\nu}\,,\quad u=\delta\,,$ (29)
@@ 4 Formulation of the problem
$\hat{F}_{uv}=-\frac{1}{2}\int_{-\infty}^{v}\partial_{i}\hat{F}_{vi}~dv^{\prime}\,.$ (30)
@@
$\Box\hat{F}_{vi}=0\,,\quad u>0\,,$ (32)
$\hat{F}_{vi}={\cal L}_{\zeta}\bar{F}_{vi}=\partial_{v}(f\bar{F}_{vi})\,,\quad u=0\,.$ (33)
@@
It should be emphasized that if there are constraints in the model, the data in the Cauchy problem cannot be arbitrary, but must also satisfy these constraints. In the case of electrodynamics, this is fulfilled, which is easy to verify by comparing (29) with (30), (31). Taking into account the constraints when setting the Cauchy problem is a substantial difference between the gauge model and the scalar model considered in Fursaev:2024czx .
We solve problem (32), (33) for transition radiation when $\bar{F}_{\mu\nu}$ is EM field of compact source (a magnetar). It is convenient to choose a frame in which the magnetar is at rest at the origin of coordinates after crossing the shockwave front. This means that in the chosen frame, before crossing the front, the magnetar was moving with some velocity. As was shown in Fursaev:2024czx , when calculating the data on $u=0$ in the linear order, one can equivalently consider the field of the source located at the origin before crossing the front, but redefine the shockwave profile function
$f(x)\to f(x)-f(0)-x^{a}\partial_{a}f(0)\,.$ (34)
We will imply that this redefinition in (33) is done.
To find the leading effect in the transition radiation of the magnetar on GSW we approximate the field of the magnetar by the field of a magnetic dipole at the center of coordinates. The corresponding current is
$\bar{j}_{\mu}(x)=\delta_{\mu}^{i}\varepsilon_{ijk}M_{j}\,\partial_{k}\delta^{(3)}(\vec{x})\,,$ (35)
where $M_{i}$ is a magnetic moment (here Latin indices correspond to spatial coordinates $x,y_{1},y_{2}$). Such a current creates a magnetic field with vector-potential
$\bar{A}_{\mu}(x)=\delta_{\mu}^{i}\varepsilon_{ijk}M_{j}\partial_{k}\phi\,,\quad\phi=-1/(4\pi r)\,,\quad r^{2}=x^{2}+y_{i}^{2}\,,$ (36)
see Landau2 . The Cauchy data (33) for the transition radiation are determined by components
$\bar{F}_{vi}(x)=(\partial_{v}\mathcal{D}_{i}^{b}-\partial_{i}\mathcal{D}_{v}^{b})\partial_{b}\phi(x)\,,$ (37)
where $i$ corresponds to $y_{1}=y,\,y_{2}=z$ and for brevity the following notations are used:
$2\mathcal{D}_{v}^{a}=M_{y}\delta_{z}^{a}-M_{z}\delta_{y}^{a}\,,\quad\mathcal{D}_{y}^{a}=-M_{x}\delta_{z}^{a}+2M_{z}\delta_{v}^{a}\,,\quad\mathcal{D}_{z}^{a}=M_{x}\delta_{y}^{a}-2M_{y}\delta_{v}^{a}\,.$ (38)
It is convenient to use coordinates $(U,r,\Omega)$, where $U=t-r$ and $\Omega$ is a solid angle which parametrizes direction of the unit vector $\vec{e}(\Omega)=(x/r,y^{i}/r)$. The hypersurface $U=0$, $t>0$ is the future light cone of the event when the shockwave hits the magnetar. All time-like trajectories end up inside the cone, $U>0$. By following a method of Fursaev:2024czx one can show that distant observers measure the following large $r$ asymptotic at $U>0$
$\hat{F}_{\mu\nu}(U,r,\Omega)\simeq\frac{1}{r}a_{\mu\nu}(U,\Omega)+O(r^{-2})\,,$ (39)
5 Transition radiation on GSW produced by ultrarelativistic objects and Fast Radio Bursts
As a model of a compact ultrarelativistic object we take a massless point-like particle which moves along the $x$-axis and has the profile function
$f(x)=8GE_{p}\log\sqrt{(y+a)^{2}+z^{2}}\,,$ (40)
where $E_{p}$ is the energy of the particle and $a$ is the impact parameter between the particle and the magnetar.
Solution to Cauchy problem (32), (33) for this profile allows us to find asymptotic (39) and the angular distribution of the radiation flux
$I(U,\Omega)=r^{2}T_{U}^{r}(U,r,\Omega)\,,$ (41)
where $T_{U}^{r}$ are components of the stress-energy tensor of the transition radiation in coordinates $U,r,\Omega$. The results are presented on the figure 2.